OpenStudy (anonymous):

The Nutty Shoppe has 10 kg of mixed pecans and cashews which sell for $8.40 per kilogram. Cashews alone sell for $8 per kilogram, and pecans sell for $9 per kilogram. How many kilograms of pecans are there in the mix?@ßαbчღ @Nurali

OpenStudy (nurali):

The total mixture sells for 10 x $8.40 = $84 Let the weight of cashews in the mix = C Then the weight of pecans in the mix = 10 - C 8C + 9(10 - C) = 84 90 - C = 84 C = 6 There are 6 kilograms of cashews in the mix.